OpenStudy (anonymous):

A set of equations is given below: Equation A: e = 4f + 1 Equation B: e = 3f + 5 Which of the following is a step that can be used to find the solution to the set of equations? 4f + 1= 3f 4f = 3f + 5 4f + 5 = 3f + 1 4f + 1 = 3f + 5

OpenStudy (anonymous):

@Luigi0210 plz

OpenStudy (anonymous):

@BOOKER23 @Ilovecake @domebotnos

OpenStudy (luigi0210):

Since e=4f+1 and e also =3f+5, you can set the equations equal to one another: \(4f+1=3f+5\)

OpenStudy (anonymous):

can you help with a few more @Luigi0210

OpenStudy (anonymous):

A set of equations is given below: Equation C: y = 4x + 8 Equation D: y = 4x + 2 Which of the following best describes the solution to the given set of equations? No solution One solution Two solutions Infinitely many solutions

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@robtobey can u help ?

OpenStudy (luigi0210):

if you do the first few steps and try and cancel out a variable, you would see that you get 0=-6 or something.. which is no solution.

OpenStudy (luigi0210):

No problem (✿◠‿◠)

OpenStudy (anonymous):

@luigi0210 Two lines, C and D, are represented by the equations given below: Line C: y = x + 14 Line D: y = 3x + 2 Which of the following shows the solution to the system of equations and explains why? (6, 20), because both lines pass through this point (6, 20), because the point does not lie on any axis (3, 11), because one of the lines passes through this point (3, 11), because the point lies between the two axes

OpenStudy (anonymous):

@Nnesha

OpenStudy (luigi0210):

Sorry about that, OS is.. not the best. Anyways, you should do the same thing and set them equal to each other: \(x+14=3x+2\) Get x to the right and numbers to the left: \(12=2x\) Now get x: \(6=x\) Now just plug in 6 and solve: y=(6)+14 Solution: \((x, y) ===> (6, --) \)

OpenStudy (anonymous):

@Luigi0210 Two lines, A and B, are represented by the equations given below: Line A: x + y = 2 Line B: 2x + y = 4 Which statement is true about the solution to the set of equations? There are infinitely many solutions. There are two solutions. There is one solution. There is no solution.

OpenStudy (luigi0210):

You have a lot of questions xD Anyways, solve it: 2x+(-x+2)=4 x=2 2+y=2 y=0 So it has one solution :P